Year 1 Year 2 Year 1 Year 2 Year 2 Revenue 3029 176896 238242 191901 1838 144924 COGS Gross Profit EBITDA 1191 37 5674 3442 212. What is company 2s current ratio in Year 2? 13. What is company 3s net margin in Year 2? 14. What is company 2s invent
Answer #1

12) Calculation of company'2 current ratio in year 2:

Current ratio= current assets/current liabilities

Current ratio= 9439/3035= 3.11

13) Calculation of company 3 net margin in year 2:

Net margin= net income/sales*100

Net margin= 3915/238242*100= 1.64%

14) Calculation of inventory Turnover of company 2 in year 2

Inventory Turnover= cost of goods sold/average inventory

Average inventory= (10271+534)/2= 5402.50

Inventory Turnover= 87544/5402.5= 16.20

15) calculation of company 3 long term debt to asset ratio for year 1:

Long term debt to asset ratio= long term debt/total assets

Long term debt to asset ratio= 69215/244587= 0.28